I need to find the length of one side of a triangle. If a triangle that has the second side 10 feet shorter than the first side. The third side is 50 feet longer than the first side. The perimeter is 684 feet. How long is the first side?
Answer by checkley77(12844)

S=F-10
T=F+50
F=FIRST SIDE
F+(F-10)+(F+50)=684
3F-10+50=684
3F=684-50+10
3F=644
F=644/3
F=214 2/3 IS THE LENGTH OF THE FIRST SIDE.
214 2/3-10=204 2/3 FOR THE SECOND SIDE.
214 2/3+50=264 2/3 FOR THE THIRD SIDE.
pROOF:
214 2/3+204 2/3+264 2/3=684
682+6/3=684
684=684